.He said the banking system could lose deposits in several ways. One channel in recent quarters has been domestic fund managers allocating away from cash and deposits to other asset classes - particularly overseas assets, Mr Jolly said Source: Banks face cash squeeze as fund managers, households draw deposits (AFR, 15.07.2018) This quote contains some misleading statements. The following questions will help you to get a deeper understanding. These tasks are also exam relevant. Assets ESF OZ Bank Liabilities and equity AUD 810m AUD 150m Deposits Loans AUD 850m Deposits Fund Manager AUD 100m Household AUD 50m Equity AUD 40m Assets Deposits at OZ Bank Fund Manager Liabilities and equity AUD 100m AUD 100m Equity Assets ESF Loans Bank B Liabilities and equity AUD 700m AUD 100m AUD 100m Deposits AUD 900m Deposits Mr Red. Equity AUD 200m a) Fund manager buys AUD 100m bonds from Mr. Red using his deposits. Mr Red has his deposit account at Bank B. Show all resulting changes on the balance sheet of OZ Bank, Fund Manager, and Bank B What happens to the ratio of loans to deposits in the banking system? b) Mr. Red uses his newly acquired deposits to repay his loan at Bank B. Show the impact on Bank Bs balance sheet. What happens to the ratio of loans to deposits in the banking system? c) OZ Bank is issuing AUD 50m bonds and the buyers happen to be OZ bank customers. Add the changes to OZ Banks balance sheet.
Answer #1

a) The OZ bank needs to honor deposits of fund manager using ESF. So, it would pass following journal entry

Deposit Fund Manager   AUD100m

         ESF                                             AUD 100m               

So, the revised Balance is OZ bank         

Assets

AUD

Liabilities and Equity

AUD

ESF

50m

Deposits

810m

Loans

850m

Household

50m

Equity

40m

For Bank OZ the Ratio of loans to Deposits have increased from 850/960=0.89 to 850/860=0.99 because of reduction in deposits of the bank

When the fund manager buys bonds of Mr. Red using the Deposits at OZ bank, he will pass the following journal entry

Bonds Mr. Red            AUD 100m

   Deposits at OZ bank                          AUD 100m

So, the revised Balance is fund manager

Assets

AUD

Liabilities and Equity

AUD

Bonds Mr.Red

100m

Equity

100m

Mr. Red deposits these funds at Bank B. So for Bank B, the entry will be

ESF                AUD100m

      Deposits Mr. Red           AUD100m

So, the revised Balance Bank B

Assets

AUD

Liabilities and Equity

AUD

ESF

200m

Deposits

700m

Loans

900m

Deposits Mr. Red

200m

Equity

200m

For Bank B, the ratio of loans to deposits have decreased from 900/800=1.125 to 900/900=1 because of increase in deposits of the bank.

b) When Mr Red uses the deposits to repay his loan at Bank B , The deposits Mr. Red would decrease to 100m and Loans Assets would decrease to 800m

c) When OZ bank issued 50m bonds, it would add 50m to ESF and add Bonds in liabilities

Revised Balance sheet

So, the revised Balance is OZ bank         

Assets

AUD

Liabilities and Equity

AUD

ESF

100m

Deposits

810m

Loans

850m

Household

50m

Bonds

50m

Equity

40m